OpenStudy (anonymous):

The value V(t) of a $18,000 car after t years that depreciates at 20% per year is given by the formula V(t) = V0(b)t. What is the value of V0 and b, and what is the car's value after 5 years (rounded to the nearest cent)? Hint: Use the equation b = 1 − r to determine depreciation. V0 = $18,000, b = 0.80, and the value after 5 years is $5,898.24 V0 = $18,000, b = 0.20, and the value after 5 years is $5.76 V0 = $18,000, b = 1.20, and the value after 5 years is $44,789.76 V0 = $18,000, b = 0.80, and the value after 5 years is $5,760.00

OpenStudy (anonymous):

@GreenCat are you able to help me with this?

OpenStudy (greencat):

no sorry

OpenStudy (phi):

Hint: Use the equation b = 1 − r that is not much of a hint, unless you know what "r" is supposed to be. Do you know what they mean by r ?

OpenStudy (anonymous):

It means doubling or 2 in this case. @phi

OpenStudy (phi):

it can't mean that. they say depreciates at 20% per year depreciate means "go down in value" I guess they mean r is 0.2 (20% written as a decimal) so "b" in your formula , using the hint is 1-0.2

OpenStudy (anonymous):

Ok that makes sense. This lesson has been really hard for me to understand. Is there anyway that you might be able to talk me through this question? @phi

OpenStudy (phi):

If you have time

OpenStudy (anonymous):

Yes I do. I'm pretty sure that it's either a or d. Because b= 1-0.2 = 0.8

OpenStudy (phi):

yes.

OpenStudy (phi):

First step Do you know what 10% of 100 is ? i.e. do you know how to find 10% of 100 ?

OpenStudy (anonymous):

It would be 10/100, which is 10

OpenStudy (phi):

you would write 10% as 10/100 (which simplifies to 0.1 as a decimal) then you do 0.1 * 100 you get 10 can you do 10% of 50 ?

OpenStudy (anonymous):

It's 5

OpenStudy (phi):

yes. But I want to make sure you get the idea say you want 1/2 of 20 you would multiply 1/2 * 20 and get 10 or 1/3 of 90 you would multiply 1/3* 90 and get 30 does that make sense?

OpenStudy (phi):

percents are the same idea, except the fraction always has 100 in the bottom 20% is the fraction 20/100 and 20% of 100 you do by multiplying the fraction 20/100 * 100 you would get 20

OpenStudy (anonymous):

Yes that makes sense. I'm following you

OpenStudy (phi):

That is the first idea that you have to be comfortable with. Now the second idea 100 is reduced by 10% that is short-hand (i.e. they leave some words out) for 100 is reduced by 10% of 100 and that means 100 minus 10% of 100 or in math 100 - 0.1*100 do you know what that is ?

OpenStudy (anonymous):

Would that be 9990?

OpenStudy (phi):

first what is 0.1*100 ?

OpenStudy (anonymous):

It's 99.9

OpenStudy (phi):

you calculator is broken what is \[ \frac{100}{10} \]

OpenStudy (anonymous):

I'm sorry I do not know how I got that. It's 10. 0.1*100 = 10

OpenStudy (phi):

yes. If you are not good with arithmetic, make a note to study it more. (Khan has lots of videos ) In the meantime, you can survive using a calculator

OpenStudy (phi):

back to the 2nd idea something costs $20 then they say, "on sale! 10% off" when they say 10% off they mean 1) find 10% of 20 can you do that ?

OpenStudy (anonymous):

10/100 *20 =2

OpenStudy (phi):

yes, something costs $20 then they say, "on sale! 10% off" when they say 10% off they mean 1) find 10% of 20, which is 2 2) take 2 off of 20, i.e. figure out 20 -2

OpenStudy (phi):

in other words, the new price will be $18 ok so far?

OpenStudy (phi):

the next idea is is to write 10% off as \[ 20- \frac{10}{100} 20 = 20\left(1-\frac{10}{100} \right) \] that uses the "distributive property" 1 - 0.1 = 0.9

OpenStudy (phi):

in other words, 10% of 20 can be written as 20 - 20*0.1 = 20(1-0.1) = 20*0.9 if you multiply 20*0.9 you get 18 (the same answer as before)

OpenStudy (anonymous):

Yes! My computer messed up for a sec. I think I actually figured out the answer. If you do the equation this way. 18,000(0.80)^5 you get 147,456/25. And if you simplify that, you get 5898.24. And that is choice A!!!

OpenStudy (anonymous):

Is that right?

OpenStudy (phi):

I'm not sure how you got 147456/25, but yes that is the correct answer on my calculator I would do 0.8^5 and get 0.32768 then multiply by 18000

OpenStudy (anonymous):

I have a few more that I need help with. Do you have time to help me?

OpenStudy (phi):

one more

OpenStudy (anonymous):

The half-life of a substance is how long it takes for half of the substance to decay or become harmless (for certain radioactive materials). The half-life of a substance is 8.6 days and there is an amount equal to 15 grams now. What is the expression for the amount A(t) that remains after t days, and what is the amount of the substance remaining (rounded to the nearest tenth) after 37 days? Hint: The exponential equation for half-life is A(t) = A0(0.5)t/H, where A(t) is the final amount remaining, A0 is the initial amount, t is time, and H is the half-life. A(t) = 15(0.5)8.6t, 0.0 gram remaining A(t) = 15(0.5)t/8.6, 0.8 gram remaining A(t) = 8.6(15)(0.5)t, 0.0 gram remaining A(t) = 15(0.5)8.6/t, 12.8 grams remaining

OpenStudy (phi):

Hint: The exponential equation for half-life is A(t) = A0(0.5)t/H, where A(t) is the final amount remaining, A0 is the initial amount, t is time, and H is the half-life to use their formula you have to interpret the info they gave you. For example, you need to know A0 , which is the "initial amount" any idea what the initial amount is ?

OpenStudy (anonymous):

I think that it is 15. Is that right?

OpenStudy (phi):

yes. they say there is an amount equal to 15 grams now. so you start with 15 grams (initial means starting or first. Like initial is the first letter of your name)

OpenStudy (phi):

so put in 15 into the equation \[ A(t) = A0(0.5)^{\frac{t}{H}} \\ A(t) = 15(0.5)^{\frac{t}{H}} \]

OpenStudy (phi):

you also need H is the half-life what is H ?

OpenStudy (anonymous):

the half life would be 8.6?

OpenStudy (phi):

so we put that in \[ A(t) = 15(0.5)^{\frac{t}{8.6}} \]

OpenStudy (anonymous):

then 37 would be t?

OpenStudy (phi):

they say "t" is time (starting from now) in days so put in 37 for t \[ A(t) = 15(0.5)^{\frac{37}{8.6}} \]

OpenStudy (phi):

I guess I should write A(37) =

OpenStudy (anonymous):

I got 0.7602601167 and then when I round to the nearest tenth it would be 0.8. Right?

OpenStudy (phi):

yes. But only one of the choices has the correct formula, so we did not need to find the actual number

OpenStudy (anonymous):

So the correct answer would be B?

OpenStudy (phi):

yes

OpenStudy (anonymous):

Thank you so much! It makes so much more sense! I really appreciate it. I only have one more question now. You really helped me :) Thank you